# Increasing Batch Size Improves Convergence of Stochastic Gradient Descent with Momentum

### v1

@@ 4 Numerical Results
We examined training ResNet-18 on the CIFAR-100 dataset using not only NSHB and SHB but also baseline optimizers: SGD, Adam, AdamW, and RMSprop with constant and increasing batch sizes (see Appendix A.4 for training ResNet-18 on Tiny ImageNet). We used a computer equipped with NVIDIA A100 80GB and Dual Intel Xeon Silver 4316 2.30GHz, 40 Cores (20 cores per CPU, 2 CPUs). The software environment was Python 3.8.2, PyTorch 2.2.2+cu118, and CUDA 12.2. We set the total number of epochs to $E=200$ and the constant momentum weight as the default values in PyTorch. The learning rate was set for Adam and AdamW to $10^{-3}$, for RMSprop to $10^{-2}$, and for SGD, SHB, and NSHB to $10^{-1}$; see also Figure 1(a).
@@
Let us first consider the learning rate and batch size scheduler in Figure 1(a) with a constant batch size ($b=2^{7}$). Figure 1(b) compares the full gradient norm $\min_{e\in[E]}\|\nabla f(\bm{\theta}_{e})\|$ for training for each optimizer and indicates that SHB decreased the full gradient norm quickly. Figures 1(c) and (d) compare the empirical loss $f(\bm{\theta}_{e})$ and the test accuracy score. These figures indicate that SGD, SHB, and NSHB minimized $f$ quickly and had test accuracies of approximately 70 %. Next, let us compare Figure 1 with Figure 2 under the scheduler with the same learning rates in Figure 1(a) and doubly batch size every 20 epochs with the initial batch size $b_{0}=2^{3}$. Figures 2(b) and (c) both show that using a doubly increasing batch size results in a faster decrease of $\min_{e\in[E]}\|\nabla f(\bm{\theta}_{e})\|$ and $f(\bm{\theta}_{e})$, compared to using a constant batch size as in Figures 1(b) and (c). The numerical results in Figures 1(b) and 2(b) are supported theoretically by Theorems 3.13.4 indicating that NSHB and SHB with increasing batch sizes minimize the gradient norm of $f$ faster than with constant batch sizes. In Figures 1(d) and 2(d), it can be seen that using a doubly increasing batch size leads to improved test accuracy for all optimizers except SHB, compared to using a constant batch size. Earlier, we observed that, with a constant batch size, convergence is slower, and accuracy improvement is more gradual. On the other hand, these results suggest that using an increasing batch size leads to a faster convergence and more efficient training. Additionally, when using an increasing batch size, the optimizer’s performance is better overall, particularly in terms of having faster convergence.
@@
Now, let us compare Figure 2 ($\delta=2$) with Figure 3 ($\delta=4$) under the scheduler with the same learning rates as in Figure 1(a) and the batch size quadruply increasing every 40 epochs with an initial batch size $b_{0}=2^{3}$. From Figures 3(b) and (c), it can be observed that the larger the batch size is, the faster the decrease of the full gradient norm $\|\nabla f(\bm{\theta}_{e})\|$ and the empirical loss $f(\bm{\theta}_{e})$ are. Specifically, the quadruply increasing batch size ($\delta=4$; Figure 3) decreases the full gradient norm $\|\nabla f(\bm{\theta}_{e})\|$ and the empirical loss $f(\bm{\theta}_{e})$ more rapidly than the doubly increasing batch size ($\delta=2$; Figure 2). Figures 2(d) and 3(d) indicate that SGD and NSHB had greater than 70 % test accuracies, which implies in turn that, for SGD and NSHB, using an increasing batch size would improve generalization more than using a constant batch size (Figure 1(d)).
@@ 4.1 Discussion and future work
Fast convergence of Adam: A particularly interesting result in Figures 23 is that an increasing batch size is applicable for Adam in the sense of minimizing the full gradient norm of $f$ fastest. Hence, we can expect that Adam with an increasing batch size has a convergence rate better than the $O(\frac{1}{\sqrt{T}})$ convergence rate of NSHB and SHB in Theorems 3.3 and 3.4. In the future, we should verify the result holds theoretically.
@@
Full gradient norm and training loss versus test accuracy: As promised in Theorems 3.1 and 3.3, NSHB with increasing batch sizes ($\delta=2,4$) minimized the full gradient norm of $f$ faster than with a constant batch size (Figures 1(b), 2(b), and 3(b)). As a result, NSHB with an increasing batch size ($\delta=2,4$) minimized the training loss $f$ (Figures 1(c), 2(c), and 3(c)) and had higher test accuracies than with a constant batch size (Figures 1(d), 2(d), and 3(d)). Moreover, Figures 13 indicate that AdamW had almost the same trend. Although Adam and AdamW with increasing batch sizes minimized $f$ quickly, the test accuracy of Adam was different from the test accuracy of AdamW (Figure 3(d)). Here, we have the following insights:
@@
  • An increasing batch size quickly minimizes the full gradient norm of the training loss in both theory and practice. In particular, SGDM with an increasing batch size converges to stationary points of the training loss, as promised in our theoretical results.
@@
  • Optimal increasing batch size schedulers with which optimizers have high test accuracies should be discussed. Specifically, we need to discuss optimal $E_{m}$ and $\delta$ such that SGDM and adaptive methods (e.g., Adam and AdamW) improve generalization.
@@ 5 Conclusion
This paper presented convergence analyses of mini-batch SGDM with a constant learning rate and momentum weight. Using a constant batch size does not lead to convergence of mini-batch SGDM to stationary points of the training loss, but using an increasing batch size does lead to its convergence. This paper also provided numerical results to support our convergence analyses. In particular, using a quadruply increasing batch size had faster convergence of mini-batch SGDM than using a doubly increasing batch size. Moreover, the numerical results indicated that using an increasing batch size is also applicable for adaptive methods, such as Adam and AdamW, in the sense of minimizing the full gradient norm of the training loss. Hence, in the future, we need to verify theoretically whether adaptive methods with increasing batch sizes have faster convergence than SGDM.
(a) Learning rate and batch size versus epochs
(b) Full gradient norm versus epochs
(c) Empirical loss versus epochs
(d) Test accuracy score versus epochs
Figure 1: (a) Schedulers for each optimizer with constant learning rates and constant batch size, (b) Full gradient norm of empirical loss for training, (c) Empirical loss value for training, and (d) Accuracy score for test to train ResNet-18 on CIFAR-100 dataset.
(a) Learning rate and batch size versus epochs
(b) Full gradient norm versus epochs
(c) Empirical loss versus epochs
(d) Test accuracy score versus epochs
Figure 2: (a) Schedulers for each optimizer with constant learning rate and batch size doubly increasing every 20 epochs, (b) Full gradient norm of empirical loss for training, (c) Empirical loss value for training, and (d) Accuracy score for test to train ResNet-18 on CIFAR-100 dataset.
(a) Learning rate and batch size versus epochs
(b) Full gradient norm versus epochs
(c) Empirical loss versus epochs
(d) Test accuracy score versus epochs
Figure 3: (a) Schedulers for each optimizer with constant learning rates and batch size quadruply increasing every 40 epochs, (b) Full gradient norm of empirical loss for training, (c) Empirical loss value for training, and (d) Accuracy score for test to train ResNet-18 on CIFAR-100 dataset.
@@ A.1 Proposition and Lemma
Let $t\in\mathbb{N}$, $\bm{\xi}_{t}$ be a random variable independent of $\bm{\xi}_{j}$ ($j\in[0:t-1]$), $\bm{\theta}_{t}\in\mathbb{R}^{d}$ be independent of $\bm{\xi}_{t}$, and $\nabla f_{B_{t}}(\bm{\theta}_{t})$ be the mini-batch gradient, where $f_{\xi_{t,i}}$ ($i\in[b_{t}]$) is the stochastic gradient (see Assumption 2.1(A2)). Then, the following hold:
@@
Assumption 2.1(A3) and the independence of $b_{t}$ and $\bm{\xi}_{t}$ ensure that
@@
which, together with Assumption 2.1(A2)(i) and the independence of $\bm{\xi}_{t}$ and $\bm{\xi}_{t-1}$, implies that
@@
Assumption 2.1(A3), the independence of $b_{t}$ and $\bm{\xi}_{t}$, and (6) imply that
@@
From the independence of $\xi_{t,i}$ and $\xi_{t,j}$ ($i\neq j$) and Assumption 2.1(A2)(i), for all $i,j\in[b_{t}]$ such that $i\neq j$,
@@
Hence, Assumption 2.1(A2)(ii) guarantees that
@@
Motivated by Lemma 1 in (Liu et al., 2020), we prove the following lemma.
@@
Under Assumption 2.1, Algorithm 1 satisfies that, for all $t\in\{0\}\cup\mathbb{N}$,
$\mathbb{E}\left[\left\|\boldsymbol{m}_{t}-\left(1-\beta\right)\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})\right\rVert^{2}\right]\leq(1-\beta)^{2}\sigma^{2}\sum_{i=0}^{t}\frac{\beta^{2(t-i)}}{b_{i}},$
@@
$\left\|\boldsymbol{m}_{t}-(1-\beta)\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})\right\|^{2}$ $=(1-\beta)^{2}\left\|\sum_{i=0}^{t}\beta^{t-i}(\nabla f_{B_{i}}(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{i}))\right\|^{2}$
@@
A similar argument as in the case where $j<i$ ensures the above equation also holds for $i<j$. Hence, Proposition A.1 guarantees that, for all $t\in\mathbb{N}$,
$\mathbb{E}\left[\left\|\boldsymbol{m}_{t}-(1-\beta)\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})\right\|^{2}\right]$ $=(1-\beta)^{2}\sum_{i=0}^{t}\mathbb{E}\left[\left\langle\beta^{t-i}(\nabla f_{B_{i}}(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{i})),\beta^{t-i}(\nabla f_{B_{i}}(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{i}))\right\rangle\right]$
@@
Suppose that Assumption 2.1 holds and $(\bm{\theta}_{t})$ is the sequence generated by Algorithm 1. We define $(\bm{z}_{t})$ for all $t\in\{0\}\cup\mathbb{N}$ as
@@
$\mathbb{E}[f(\boldsymbol{z}_{t+1})]$ $\leq\mathbb{E}[f(\boldsymbol{z}_{t})]+\eta\left[L\left\{\left(\frac{\beta}{1-\beta}\right)^{2}+\frac{3}{2}\right\}\eta-1\right]\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]+\frac{L\sigma^{2}\eta^{2}}{2}\left\{\beta^{2}\sum_{i=0}^{t-1}\frac{\beta^{2(t-1-i)}}{b_{i}}+\frac{1}{b_{t}}\right\}$
@@
The descent lemma (Beck, 2017, Lemma 5.7) ensures that, for all $t\in\{0\}\cup\mathbb{N}$,
$\mathbb{E}_{\bm{\xi}_{t}}[f(\boldsymbol{z}_{t+1})]$ $\leq f(\boldsymbol{z}_{t})+\mathbb{E}_{\bm{\xi}_{t}}[\langle\nabla f(\boldsymbol{z}_{t}),\boldsymbol{z}_{t+1}-\boldsymbol{z}_{t}\rangle]+\frac{L}{2}\mathbb{E}_{\bm{\xi}_{t}}[\|\boldsymbol{z}_{t+1}-\boldsymbol{z}_{t}\|^{2}],$
which, together with $\boldsymbol{z}_{t+1}=\boldsymbol{z}_{t}-\eta\nabla f_{B_{t}}(\bm{\theta}_{t})$ (Liu et al., 2020, Lemma 3), implies that
$\mathbb{E}_{\bm{\xi}_{t}}[f(\boldsymbol{z}_{t+1})]$ $\leq f(\boldsymbol{z}_{t})-\eta\underbrace{\mathbb{E}_{\bm{\xi}_{t}}[\langle\nabla f(\boldsymbol{z}_{t}),\nabla f_{B_{t}}(\bm{\theta}_{t})\rangle]}_{X_{t}}+\frac{L\eta^{2}}{2}\underbrace{\mathbb{E}_{\bm{\xi}_{t}}[\|\nabla f_{B_{t}}(\bm{\theta}_{t})\|^{2}]}_{Y_{t}}.$ (7)
@@
$X_{t}=\langle\nabla f(\boldsymbol{z}_{t}),\mathbb{E}_{\bm{\xi}_{t}}[\nabla f_{B_{t}}(\bm{\theta}_{t})]\rangle=\langle\nabla f(\boldsymbol{z}_{t}),\nabla f(\bm{\theta}_{t})\rangle,$
@@
$-\eta X_{t}$ $=\langle\nabla f(\boldsymbol{z}_{t}),-\eta\nabla f(\bm{\theta}_{t})\rangle$
$=\langle\nabla f(\boldsymbol{z}_{t})-\nabla f(\bm{\theta}_{t}),-\eta\nabla f(\bm{\theta}_{t})\rangle-\eta\|\nabla f(\bm{\theta}_{t})\|^{2}$
$\leq(\sqrt{\eta}\|\nabla f(\boldsymbol{z}_{t})-\nabla f(\bm{\theta}_{t})\|)(\sqrt{\eta}\|\nabla f(\bm{\theta}_{t})\|)-\eta\|\nabla f(\bm{\theta}_{t})\|^{2}$
$\leq\frac{\rho\eta}{2}\|\nabla f(\boldsymbol{z}_{t})-\nabla f(\boldsymbol{\theta}_{t})\|^{2}+\frac{\eta}{2\rho}\|\nabla f(\bm{\theta}_{t})\|^{2}-\eta\|\nabla f(\bm{\theta}_{t})\|^{2}$
$\leq\frac{\rho\eta L^{2}}{2}\|\boldsymbol{z}_{t}-\boldsymbol{\theta}_{t}\|^{2}+\eta\left(\frac{1}{2\rho}-1\right)\|\nabla f(\bm{\theta}_{t})\|^{2}.$
The definitions of $\bm{z}_{t}$ and $\bm{\theta}_{t}$ ($=\bm{\theta}_{t-1}-\eta\bm{m}_{t-1}$) ensure that, for all $t\geq 1$,
$\boldsymbol{z}_{t}=\frac{1}{1-\beta}\boldsymbol{\theta}_{t}-\frac{\beta}{1-\beta}(\boldsymbol{\theta}_{t}+\eta\boldsymbol{m}_{t-1})=\boldsymbol{\theta}_{t}-\frac{\beta}{1-\beta}\eta\boldsymbol{m}_{t-1}.$
@@
$\boldsymbol{z}_{t}=\boldsymbol{\theta}_{t}-\frac{\beta}{1-\beta}\eta\boldsymbol{m}_{t-1}.$
@@
$-\eta X_{t}\leq\frac{\rho\eta^{3}L^{2}}{2}\left(\frac{\beta}{1-\beta}\right)^{2}\|\boldsymbol{m}_{t-1}\|^{2}+\eta\left(\frac{1}{2\rho}-1\right)\|\nabla f(\bm{\theta}_{t})\|^{2}.$ (8)
@@
$\|\boldsymbol{m}_{t-1}\rVert^{2}$ $\leq 2\left\|\boldsymbol{m}_{t-1}-(1-\beta)\sum_{i=0}^{t-1}\beta^{t-1-i}\nabla f(\bm{\theta}_{i})\right\rVert^{2}+2\underbrace{\left\|(1-\beta)\sum_{i=0}^{t-1}\beta^{t-1-i}\nabla f(\bm{\theta}_{i})\right\rVert^{2}}_{Z_{t}}.$ (9)

### v2

@@ 4 Numerical Results
We examined training ResNet-18 on the CIFAR-100 and Tiny-ImageNet datasets using not only NSHB and SHB but also baseline optimizers: SGD, Adam, AdamW, and RMSprop with constant and increasing batch sizes. The results on Tiny-ImageNet are provided in Appendix A.4. We used a computer equipped with NVIDIA A100 80GB and Dual Intel Xeon Silver 4316 2.30GHz, 40 Cores (20 cores per CPU, 2 CPUs). The software environment was Python 3.8.2, PyTorch 2.2.2+cu118, and CUDA 12.2. We set the total number of epochs to $E=200$ and the constant momentum weight as the default values in PyTorch. The learning rate for Adam and AdamW was set to $10^{-3}$, for RMSprop to $10^{-2}$, and for SGD, SHB, and NSHB to $10^{-1}$; see also Figure 4(a). We should note that the learning rates used in our experiments appear to be theoretically justified within the range implied by Theorem 3.2 (see Appendix A.7 for further details). All results are averaged over three independent trials, and the mean, maximum, and minimum at each epoch are shown.
Figure 4: (a) Schedules for each optimizer with constant learning rates and a constant batch size, (b) Full gradient norm of empirical loss for training, (c) Empirical loss value for training, and (d) Accuracy score for test to train ResNet-18 on CIFAR-100 dataset.
Figure 8: (a) Schedules for each optimizer with constant learning rate and a batch size doubling every 20 epochs, (b) Full gradient norm of empirical loss for training, (c) Empirical loss value for training, and (d) Accuracy score for test to train ResNet-18 on CIFAR-100 dataset.
Figure 12: (a) Schedules for each optimizer with constant learning rates and a batch size quadrupling every 40 epochs, (b) Full gradient norm of empirical loss for training, (c) Empirical loss value for training, and (d) Accuracy score for test to train ResNet-18 on CIFAR-100 dataset.
@@
Let us first consider the learning rate and batch size scheduler in Figure 4(a) with a constant batch size ($b=2^{7}$). Figure 4(b) compares the full gradient norm $\min_{e\in[E]}\|\nabla f(\bm{\theta}_{e})\|$ for training for each optimizer and indicates that SHB decreased the full gradient norm quickly. Figures 4(c) and (d) compare the empirical loss $f(\bm{\theta}_{e})$ and the test accuracy score. These figures indicate that SGD, SHB, and NSHB minimized $f$ quickly and had test accuracies of approximately 70 %. Next, let us compare Figure 4 with Figure 4 for when the scheduler uses the same learning rates as in Figure 4(a) and a batch size doubling every 20 epochs with the initial batch size set at $b_{0}=2^{3}$. Figures 4(b) and (c) both show that using a doubly increasing batch size results in a faster decrease in $\min_{e\in[E]}\|\nabla f(\bm{\theta}_{e})\|$ and $f(\bm{\theta}_{e})$, compared with using a constant batch size as in Figures 4(b) and (c). The numerical results in Figures 4(b) and 4(b) are supported theoretically by Theorems 3.1, 3.2, A.1 and A.2 indicating that NSHB and SHB with increasing batch sizes minimize the gradient norm of $f$ faster than with constant batch sizes. In Figures 4(d) and 4(d), it can be seen that using a doubly increasing batch size leads to improved test accuracy for all optimizers except SHB, compared with using a constant batch size. Earlier, we observed that, with a constant batch size, convergence is slower and accuracy improves more gradually. On the other hand, these results suggest that using an increasing batch size leads to faster convergence and more efficient training. Additionally, the optimizer’s performance is better overall when using an increasing batch size.
@@
Now, let us compare Figure 4 ($\delta=2$) with Figure 4 ($\delta=4$) when the scheduler uses the same learning rates as in Figure 4(a) and a batch size quadrupling every 40 epochs with the initial batch size set at $b_{0}=2^{3}$. From Figures 4(b) and (c), it can be observed that the larger the batch size is, the faster the decrease of the full gradient norm $\|\nabla f(\bm{\theta}_{e})\|$ and the empirical loss $f(\bm{\theta}_{e})$ become. Specifically, the quadruply increasing batch size ($\delta=4$; Figure 4) decreases the full gradient norm $\|\nabla f(\bm{\theta}_{e})\|$ and the empirical loss $f(\bm{\theta}_{e})$ more rapidly than the doubly increasing batch size ($\delta=2$; Figure 4). Figures 4(d) and 4(d) indicate that SGD and NSHB had test accuracies greater than 70 %, which implies that, for SGD and NSHB, using an increasing batch size would improve generalization more than using a constant batch size (Figure 4(d)).
@@ 4.1 Discussion and future work
Fast convergence of Adam: A particularly interesting result in Figures 44 is that an increasing batch size is applicable for Adam in the sense of it helping to minimize the full gradient norm of $f$ fastest. Hence, we can expect that Adam with an increasing batch size has a convergence rate better than the $O(\frac{1}{\sqrt{T}})$ convergence rate of NSHB and SHB in Theorems 3.2 and A.2. In the future, we should verify that this result holds theoretically.
@@
Full gradient norm and training loss versus test accuracy: As promised in Theorems 3.1 and 3.2, NSHB with increasing batch sizes ($\delta=2,4$) minimized the full gradient norm of $f$ faster than with a constant batch size (Figures 4(b), 4(b), and 4(b)). As a result, NSHB with an increasing batch size ($\delta=2,4$) minimized the training loss $f$ (Figures 4(c), 4(c), and 4(c)) and had higher test accuracies than with a constant batch size (Figures 4(d), 4(d), and 4(d)). Moreover, Figures 44 indicate that AdamW had almost the same trend. Although Adam and AdamW with increasing batch sizes both minimized $f$ quickly, their test accuracies were different (Figure 4(d)). Here, we have the following insights:
@@
(1) An increasing batch size quickly minimizes the full gradient norm of the training loss in both theory and practice. In particular, SGDM with an increasing batch size converges to stationary points of the training loss, as promised in our theoretical results.
@@
(2) Optimal increasing-batch size-schedulers with which optimizers have high test accuracies should be discussed. Specifically, we need to find the optimal $E_{m}$ and $\delta$ with which SGDM and adaptive methods (e.g., Adam and AdamW) can improve generalization.
4.2 Computational cost evaluation
We evaluated the efficiency of fixed and increasing batch-size schedules in terms of the number of stochastic gradient computations required to achieve specific training goals. To quantify this, we define the SFO complexity, which corresponds to the total number of gradient evaluations. If the batch size is $b$ and the number of training steps is $T$, then the SFO complexity is given by $Tb$.
To simulate realistic GPU memory constraints, we capped the maximum batch size at 1024. All experiments were conducted using the CIFAR-100 dataset with the NSHB optimizer, under the same settings as in the other numerical experiments.
We compared the SFO complexity required to (i) reach a gradient norm threshold (e.g., $\|\nabla f(\bm{\theta}_{t})\|<0.05$) and (ii) achieve 70% test accuracy (see also Appendices A.5 and A.6).
Experimental settings. We considered both fixed and increasing batch size schedules in the gradient norm and test accuracy evaluations. For the gradient norm evaluation, the fixed setting used $b=8$ and $b=128$, while the increasing setting started with $b=8$ (doubling every 20 epochs) and with $b=128$ (doubling every 50 epochs). For the test accuracy evaluation, the fixed setting also used $b=8$ and $b=128$, while the increasing setting started with $b=8$ (doubling every 20 epochs) and with $b=128$ (doubling every 25 epochs).
Results.
Table 2: SFO complexity to reach gradient norm threshold ($b=8$)
Method $\|\nabla f(\bm{\theta}_{t})\|<0.1$ $\|\nabla f(\bm{\theta}_{t})\|<0.05$
Fixed batch size ($b=8$) 2,750,000 5,250,000
Increasing batch size (initial $b=8$) 2,050,016 2,500,160
Table 3: SFO complexity to reach gradient norm threshold ($b=128$)
Method $\|\nabla f(\bm{\theta}_{t})\|<0.1$ $\|\nabla f(\bm{\theta}_{t})\|<0.06$
Fixed batch size ($b=128$) 5,755,520 9,809,408
Increasing batch size (initial $b=128$) 2,903,808 5,061,376
Table 4: SFO complexity to reach 70% test accuracy
Method SFO
Fixed batch size ($b=8$) 5,250,000
Increasing batch size (initial $b=8$) 2,050,016
Fixed batch size ($b=128$) 2,502,400
Increasing batch size (initial $b=128$) 1,301,376
Discussion. These results show that increasing the batch size significantly reduces the total number of stochastic gradient computations needed to achieve optimization and generalization goals, especially under realistic memory constraints. Our experiments confirm that using an increasing batch size reduces gradient evaluations compared with using a fixed batch size, even when both achieve similar test accuracy and optimization performance. This highlights that larger batch sizes are not just a theoretical convenience but offer clear practical benefits. They provide an efficient way to reduce training costs without compromising generalization, especially in large-scale deep learning under memory and compute constraints.
@@ 5 Conclusion
This paper presented convergence analyses of mini-batch SGDM with a constant learning rate and momentum weight. We showed that, unlike prior studies that assume a decaying learning rate to ensure convergence, increasing the batch size under a constant learning rate and momentum not only guarantees convergence but also achieves faster convergence. Numerical experiments supported our theory, demonstrating faster convergence, higher test accuracy, and reduced computational costs compared with a constant batch size. Moreover, our results suggested that increasing batch size can also benefit adaptive methods such as Adam and AdamW. Future work includes extending our analysis to larger-scale datasets and deeper architectures, as well as generalizing the framework beyond the exponential growth schedule to cover polynomial growth and adaptive schemes. These directions of study will further clarify the role of batch size in modern optimization and strengthen the connection between theory and practice.
Acknowledgements
We are sincerely grateful to Program Chairs, Area Chairs, and the four anonymous reviewers for helping us improve the original manuscript. This work was supported by the Japan Society for the Promotion of Science (JSPS) KAKENHI Grant Number 24K14846 awarded to Hideaki Iiduka.
@@ A.1 Proposition and lemma
Let $t\in\mathbb{N}$, $\bm{\xi}_{t}$ be a random variable independent of $\bm{\xi}_{j}$ ($j\in[0:t-1]$), $\bm{\theta}_{t}\in\mathbb{R}^{d}$ be independent of $\bm{\xi}_{t}$, and $\nabla f_{B_{t}}(\bm{\theta}_{t})$ be the mini-batch gradient, where $f_{\xi_{t,i}}$ ($i\in[b_{t}]$) is the stochastic gradient (see Assumption 1(A2)). Then, the following hold:
@@
Assumption 1(A3) and the independence of $b_{t}$ and $\bm{\xi}_{t}$ ensure that
@@
which, together with Assumption 1(A2)(i) and the independence of $\bm{\xi}_{t}$ and $\bm{\xi}_{t-1}$, implies that
@@
Assumption 1(A3), the independence of $b_{t}$ and $\bm{\xi}_{t}$, and (6) imply that
@@
From the independence of $\xi_{t,i}$ and $\xi_{t,j}$ ($i\neq j$) and Assumption 1(A2)(i), for all $i,j\in[b_{t}]$ such that $i\neq j$,
@@
Hence, Assumption 1(A2)(ii) guarantees that
@@
Motivated by Lemma 1 in [Liu et al., 2020], we prove the following lemma.
@@
Under Assumption 1, Algorithm 1 satisfies that, for all $t\in\{0\}\cup\mathbb{N}$,
$\mathbb{E}\left[\left\|\bm{m}_{t}-\left(1-\beta\right)\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})\right\rVert^{2}\right]\leq(1-\beta)^{2}\sigma^{2}\sum_{i=0}^{t}\frac{\beta^{2(t-i)}}{b_{i}},$
@@
$\left\|\bm{m}_{t}-(1-\beta)\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})\right\|^{2}$
$=(1-\beta)^{2}\left\|\sum_{i=0}^{t}\beta^{t-i}(\nabla f_{B_{i}}(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{i}))\right\|^{2}$
@@
A similar argument as in the case of $j<i$ ensures the above equation also holds for $i<j$. Hence, Proposition A.1 guarantees that, for all $t\in\mathbb{N}$,
$\mathbb{E}\left[\left\|\bm{m}_{t}-(1-\beta)\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})\right\|^{2}\right]$
$=(1-\beta)^{2}\sum_{i=0}^{t}\mathbb{E}\left[\left\langle\beta^{t-i}(\nabla f_{B_{i}}(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{i})),\beta^{t-i}(\nabla f_{B_{i}}(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{i}))\right\rangle\right]$
@@
Suppose that Assumption 1 holds and $(\bm{\theta}_{t})$ is the sequence generated by Algorithm 1. We define $(\bm{z}_{t})$ for all $t\in\{0\}\cup\mathbb{N}$ as
@@
$\mathbb{E}[f(\bm{z}_{t+1})]$ $\leq\mathbb{E}[f(\bm{z}_{t})]+\eta\left[L\left\{\left(\frac{\beta}{1-\beta}\right)^{2}+\frac{3}{2}\right\}\eta-1\right]\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]$
$\quad+\frac{L\sigma^{2}\eta^{2}}{2}\left\{\beta^{2}\sum_{i=0}^{t-1}\frac{\beta^{2(t-1-i)}}{b_{i}}+\frac{1}{b_{t}}\right\}$
@@
The descent lemma [Beck, 2017, Lemma 5.7] ensures that, for all $t\in\{0\}\cup\mathbb{N}$,
$\mathbb{E}_{\bm{\xi}_{t}}[f(\bm{z}_{t+1})]$ $\leq f(\bm{z}_{t})+\mathbb{E}_{\bm{\xi}_{t}}[\langle\nabla f(\bm{z}_{t}),\bm{z}_{t+1}-\bm{z}_{t}\rangle]+\frac{L}{2}\mathbb{E}_{\bm{\xi}_{t}}[\|\bm{z}_{t+1}-\bm{z}_{t}\|^{2}],$
which, together with $\bm{z}_{t+1}=\bm{z}_{t}-\eta\nabla f_{B_{t}}(\bm{\theta}_{t})$ [Liu et al., 2020, Lemma 3], implies that
$\mathbb{E}_{\bm{\xi}_{t}}[f(\bm{z}_{t+1})]$ $\leq f(\bm{z}_{t})-\eta\underbrace{\mathbb{E}_{\bm{\xi}_{t}}[\langle\nabla f(\bm{z}_{t}),\nabla f_{B_{t}}(\bm{\theta}_{t})\rangle]}_{X_{t}}+\frac{L\eta^{2}}{2}\underbrace{\mathbb{E}_{\bm{\xi}_{t}}[\|\nabla f_{B_{t}}(\bm{\theta}_{t})\|^{2}]}_{Y_{t}}.$ (7)
@@
$X_{t}=\langle\nabla f(\bm{z}_{t}),\mathbb{E}_{\bm{\xi}_{t}}[\nabla f_{B_{t}}(\bm{\theta}_{t})]\rangle=\langle\nabla f(\bm{z}_{t}),\nabla f(\bm{\theta}_{t})\rangle,$
@@
$-\eta X_{t}$ $=\langle\nabla f(\bm{z}_{t}),-\eta\nabla f(\bm{\theta}_{t})\rangle$
$=\langle\nabla f(\bm{z}_{t})-\nabla f(\bm{\theta}_{t}),-\eta\nabla f(\bm{\theta}_{t})\rangle-\eta\|\nabla f(\bm{\theta}_{t})\|^{2}$
$\leq(\sqrt{\eta}\|\nabla f(\bm{z}_{t})-\nabla f(\bm{\theta}_{t})\|)(\sqrt{\eta}\|\nabla f(\bm{\theta}_{t})\|)-\eta\|\nabla f(\bm{\theta}_{t})\|^{2}$
$\leq\frac{\rho\eta}{2}\|\nabla f(\bm{z}_{t})-\nabla f(\bm{\theta}_{t})\|^{2}+\frac{\eta}{2\rho}\|\nabla f(\bm{\theta}_{t})\|^{2}-\eta\|\nabla f(\bm{\theta}_{t})\|^{2}$
$\leq\frac{\rho\eta L^{2}}{2}\|\bm{z}_{t}-\bm{\theta}_{t}\|^{2}+\eta\left(\frac{1}{2\rho}-1\right)\|\nabla f(\bm{\theta}_{t})\|^{2}.$
The definitions of $\bm{z}_{t}$ and $\bm{\theta}_{t}$ ($\bm{\theta}_{t-1}-\eta\bm{m}_{t-1}$) ensure that, for all $t\geq 1$,
$\bm{z}_{t}=\frac{1}{1-\beta}\bm{\theta}_{t}-\frac{\beta}{1-\beta}(\bm{\theta}_{t}+\eta\bm{m}_{t-1})=\bm{\theta}_{t}-\frac{\beta}{1-\beta}\eta\bm{m}_{t-1}.$
@@
$\bm{z}_{t}=\bm{\theta}_{t}-\frac{\beta}{1-\beta}\eta\bm{m}_{t-1}.$
@@
$-\eta X_{t}\leq\frac{\rho\eta^{3}L^{2}}{2}\left(\frac{\beta}{1-\beta}\right)^{2}\|\bm{m}_{t-1}\|^{2}+\eta\left(\frac{1}{2\rho}-1\right)\|\nabla f(\bm{\theta}_{t})\|^{2}.$ (8)
@@
$\|\bm{m}_{t-1}\rVert^{2}$ $\leq 2\left\|\bm{m}_{t-1}-(1-\beta)\sum_{i=0}^{t-1}\beta^{t-1-i}\nabla f(\bm{\theta}_{i})\right\rVert^{2}+2\underbrace{\left\|(1-\beta)\sum_{i=0}^{t-1}\beta^{t-1-i}\nabla f(\bm{\theta}_{i})\right\rVert^{2}}_{Z_{t}}.$ (9)